### v1

@@ 1 Introduction
Large language models (LLMs) have significantly transformed the field of natural language processing, achieving remarkable performance in text generation, summarization and sentiment analysis (Brown et al., 2020; OpenAI, 2023; Dubey et al., 2024; Rivière et al., 2024; Zhang et al., 2024b; Yeo et al., 2024). Despite the impressive capabilities of LLMs, the vast majority of them still are very much English-centric (Wendler et al., 2024; Zhong et al., 2024).
Unfortunately, this situation has led LLMs in regions with many under-represented languages such as Southeast Asia (SEA) to suffer. Languages with lower resources, such as Filipino, Lao, Burmese and Khmer in the SEA region, are not supported by many open-source English-centric LLMs such as Llama (Dubey et al., 2024) and Olmo (Groeneveld et al., 2024). This raises a pressing need to mitigate this language resource and representation gap between English and SEA languages.
Recently, there have been many attempts to create multilingual LLMs in an open-source manner. For instance, BLOOM (Scao et al., 2022) was a project aimed at increasing multilingual presence in open-source LLMs by supporting 46 natural languages. Popular LLM families such as Llama (Dubey et al., 2024), Gemma (Rivière et al., 2024) and Qwen (Yang et al., 2024a) have also introduced multilingual LLMs for their latest iteration. During our evaluations, we found that the performance of these models is acceptable in the general case, i.e., if we consider evaluation benchmarks formulated from English datasets, but we observe that the performance degrades on SEA-specific benchmarks.
Moreover, researchers have also introduced LLMs such as SeaLLMs Nguyen et al. (2024); Zhang et al. (2024a) and Sailor Dou et al. (2024) to specifically address the LLM gap in SEA languages. However, the performance of these models is less than ideal for languages such as Thai or Tamil [^2] Tamil is one of the official languages in Singapore. It is also spoken in other areas in the SEA region, such as Malaysia. (10X et al., 2024; AI Products Team, 2024).
@@ 2.1 Pre-training data
The CPT data consists of a curated set of English, multilingual and code corpora from several open source repositories like Dolma (Soldaini et al., 2024), FineWeb (Penedo et al., 2024), the-stackv2 (Lozhkov et al., 2024), SEA-LION-Pile (AI Singapore, 2023), SEA-LION-Pile-v2 (AI Singapore, 2025), as well as documents from CommonCrawl (CommonCrawl, 2024) and from the public domain such as Wikipedia (Foundation, 2024). For SEA-LION-Pilev2, we filter CommonCrawl WARC data for documents in SEA languages (i.e., Burmese, Simplified Chinese, Indonesian, Khmer, Lao, Malay, Filipino, Tamil, Thai and Vietnamese) using the pretrained fasttext language classifier (Joulin et al., 2017). A document is retained if the language code reported in its metadata matches with that of one of the aforementioned SEA languages. Additionally, we further clean up the data with Trafilatura (Barbaresi, 2021).
To determine the best dataset ratio between SEA languages, code and English for the CPT process, we perform a series of small-scale CPT experiments each with a training budget of 10B tokens and varying proportions of English, code and SEA language data. We settled on an optimal data mix ratio of 55% SEA languages, 25% English and 20% code tokens for a budget of 200B tokens. For a detailed breakdown of the token count by languages, please refer to the model card. [^5] Gemma-SEA-LION-v3-9B Model Card
@@ 3.1 Post-training data
The post-training data for instruction fine-tuning consists of Infinity-Instruct [Foundation and Chat] (Beijing Academy of Artificial Intelligence, 2024), OpenMath-Instruct 2 (Toshniwal et al., 2024) and our own SEA-Instruct. In particular, SEA-Instruct consists of multiple open-source instruction datasets, a synthetically generated dataset following the Magpie (Xu et al., 2024) template, and hand-crafted datasets collected from native Southeast Asians. The full detail of the SEA-Instruct and SEA-Preference dataset can be found in the model card. [^6] Gemma-SEA-LION-v3-9B-IT Model Card
@@ 4.1 Setup
For the evaluation, we compared our models against well-known LLMs, such as SeaLLMsv3 Zhang et al. (2024a), Sailorv2 Team (2024), Qwen 2.5 Yang et al. (2024a), Gemma 2 Rivière et al. (2024) and Llama 3.1 Dubey et al. (2024), where the parameters of those models are less than 10 billion parameters, similar to our models. The evaluations are split into two areas as follows.
@@ 4.3 Performance Analysis
Merge 2: Incorporating instruct models To reintroduce helpfulness, relevance and informativeness of responses observed in Llama 3.1 and Gemma 2 models, we perform further merges of open-source instruct models. While we observed significant increases in MT-Bench benchmark scores for Vietnamese and Thai, we also observed a slight degradation of average SEA-HELM performance as well as a slight degradation of Indonesian MTBench scores, which we view as acceptable tradeoffs for the significant performance increases in Vietnamese and Thai. Note that since Merge-1-Gemma already demonstrates great performance across the SEA-HELM benchmark, we choose to skip this step for the Gemma model.
@@
The alignment step significantly pushes the model towards longer, more helpful and sensitive responses, but also compromises performance in more task-specific benchmarks and instruction following in some languages, which we try to resolve in the next step.

### v2

@@ 1 Introduction
Large language models (LLMs) have significantly transformed the field of natural language processing, achieving remarkable performance in text generation, summarization and sentiment analysis (Brown et al., 2020; OpenAI, 2023; Dubey et al., 2024; Rivière et al., 2024; Zhang et al., 2024b; Yeo et al., 2024).
Despite their impressive capabilities, most LLMs remain heavily English-centric (Wendler et al., 2024; Zhong et al., 2024). Unfortunately, this situation has led LLMs in regions with many under-represented languages such as Southeast Asia (SEA) to suffer. Languages with lower resources, such as Filipino, Lao, Burmese and Khmer in the SEA region, are not supported by many open-source English-centric LLMs. This underscores the need to bridge the resource and representation gap between English and SEA languages.
Recently, there have been many attempts to create multilingual LLMs in an open-source manner, e.g., BLOOM (Scao et al., 2022), a project aimed at increasing multilingual presence in open-source LLMs by supporting 46 natural languages. Popular LLM families such as Llama (Dubey et al., 2024), Gemma (Rivière et al., 2024) and Qwen (Yang et al., 2024a) have also introduced multilingual LLMs for their latest iteration. During our evaluations, we found that the performance of these models is acceptable in the general case, i.e., if we consider evaluation benchmarks formulated from English datasets, but we observe that the performance degrades on SEA-specific benchmarks. Moreover, researchers have also introduced LLMs such as SeaLLMs Nguyen et al. (2024); Zhang et al. (2024a) and Sailor Dou et al. (2024) to specifically address the LLM gap in SEA languages. However, the performance of these models is less than ideal for languages such as Thai or Tamil [^2] Tamil is one of the official languages in Singapore. It is also spoken in other areas in the SEA region, such as Malaysia. (10X et al., 2024; AI Products Team, 2024).
@@ 2.1 Pre-training data
The CPT data consists of a curated set of English, multilingual and code corpora from several open source repositories like Dolma (Soldaini et al., 2024), FineWeb (Penedo et al., 2024), the-stackv2 (Lozhkov et al., 2024), SEA-LION-Pile (AI Singapore, 2023), SEA-LION-Pile-v2 (AI Singapore, 2025), as well as documents from CommonCrawl (CommonCrawl, 2024) and from the public domain such as Wikipedia (Foundation, 2024). For SEA-LION-Pilev2, we filter CommonCrawl WARC data for documents in SEA languages (i.e., Burmese, Simplified Chinese, Indonesian, Khmer, Lao, Malay, Filipino, Tamil, Thai and Vietnamese) using the pretrained fasttext language classifier (Joulin et al., 2017).
A document is retained if the language code reported in its metadata matches with that of one of the aforementioned SEA languages. Additionally, we further clean up the data with Trafilatura (Barbaresi, 2021). To determine the best dataset ratio between SEA languages, code and English for the CPT process, we perform a series of small-scale CPT experiments each with a training budget of 10B tokens and varying proportions of English, code and SEA language data. We settled on an optimal data mix ratio of 55% SEA languages, 25% English and 20% code tokens for a budget of 200B tokens. For a detailed breakdown of the token count by languages, please refer to the model card. [^5] Gemma-SEA-LION-v3-9B Model Card
@@ 3.1 Post-training data
The post-training data for instruction fine-tuning consists of Infinity-Instruct [Foundation and Chat] (Beijing Academy of Artificial Intelligence, 2024), OpenMath-Instruct 2 (Toshniwal et al., 2024) and our own SEA-Instruct.
In particular, SEA-Instruct consists of multiple open-source instruction datasets, a synthetically generated dataset following the Magpie (Xu et al., 2024) template, and hand-crafted datasets collected from native Southeast Asians. The full detail of the SEA-Instruct and SEA-Preference dataset can be found in the model card. [^6] Gemma-SEA-LION-v3-9B-IT Model Card
@@ 4.1 Setup
For the evaluation, we compared our models against well-known LLMs: SeaLLMsv3 Zhang et al. (2024a), Sailorv2 Team (2024), Qwen 2.5 Yang et al. (2024a), Gemma 2 Rivière et al. (2024) and Llama 3.1 Dubey et al. (2024), where the parameters of those models are less than 10 billion parameters, similar to our models. The evaluations are split into two areas as follows.
@@ 4.3 Performance Analysis
Merge 2: Incorporating instruct models To reintroduce helpfulness, relevance and informativeness of responses observed in Llama 3.1 and Gemma 2 models, we perform further merges of open-source instruct models. While we observed significant increases in MT-Bench benchmark scores for Vietnamese and Thai, we also observed a slight degradation of average SEA-HELM performance as well as a slight degradation of Indonesian MTBench scores, which we view as acceptable tradeoffs for the significant performance increases in Vietnamese and Thai.
@@
The alignment step encourages longer, more helpful and sensitive responses but hurts performance on task-specific benchmarks and instruction following in some languages – an issue we address in the next step.